The common oxidation number for an alkaline earth metal is +2. Using the Born–Mayer equation and a Born–Haber cycle, show that CaCl is an exothermic compound. Use a suitable analogy to estimate an ionic radius for Ca+. The sublimation enthalpy of Ca(s) is 176 kJmol-1. Show that an explanation for the nonexistence of CaCl can be found in the enthalpy change for
the reaction 2 CaCl(s) → Ca(s) + CaCl2(s).
1
Expert's answer
2020-06-08T15:45:13-0400

continuation:

"=-782771\frac{J}{mol}\approx-783\frac{kJ}{mol}"

Then ΔHf="534.9\\frac{kJ}{mol}\u200b+U_{L}=534.9\\frac{kJ}{mol}\u200b-783\\frac{kJ}{mol}\u200b\\approx-248\\frac{kJ}{mol}" < 0 - an exothermic process

Thus we show that CaCl is an exothermic compound.

for the reaction 2 CaCl(s) → Ca(s) + CaCl2(s).

"\Delta H=\Delta H_{f,CaCl2}-2\cdot
\Delta H_{f,CaCl}=-785.8\frac{kJ}{mol}-2\cdot(-248\frac{kJ}{mol})\approx-289.8\frac{kJ}{mol}" - exotermic. Thus nonexistent CaCl would decompose spontaneously to CaCl2 and Ca.